The normals at any point P of the ellipsoid x
2
9 +
y
2
4 +z
2 = 1 meet the coordinate
planes in Q1,Q2,Q3, respectively. Show that PQ1 : PQ2 : PQ3 :: 9 : 4 : 1.
1
Expert's answer
2020-03-03T10:58:24-0500

As per the given condition in the question,

The equation of the ellipse,

"\dfrac{x^2}{9}+\dfrac{y^2}{4}+z^2=1"

now compare the above the general equation of the ellipse,

"\dfrac{x^2}{a^2}+\dfrac{y^2}{b^2}+\dfrac{z^2}{c^2}=1"

Let the parametric point "P(a\sin\theta\cos\phi, b\cos\theta\sin\phi, c\cos\theta)"

a=3, b=2, c=1

now putting "(a\sin\theta\cos\phi, 0, 0)",

We know from the general equation,

we know that from the general equation,

"x=-\dfrac{a}{a^2-b^2-c^2}=\dfrac{9}{-4}"

similarly "y=-\dfrac{b}{a^2-b^2-c^2}=\dfrac{4}{-4}"

similarly "z=\dfrac{c}{a^2-b^2-c^2}=\dfrac{-1}{4}"

So, distance "PQ_1=\sqrt{(x-0)^2+(0-0)^2+(0-0)^2}=\dfrac{9}{4}"

Similarly "PQ_2=\dfrac{4}{4}"

"PQ_3=\dfrac{1}{4}"

Hence the required ratio is 9:4:1